a. A car dealer believes that average daily sales for four different dealerships in four separate states are equal. A random sample of days results in the following data on daily sales:

Ohio                New York       West Virginia              Pennsylvania

               3                          10                         3                                  20

               2                            0                         4                                  11

               6                            7                         5                                  8

               4                            8                                                             2

               4                            0                                                             14

               7

               2

Use ANOVA to test this claim at the 0.05 level.

b. Use Excel to support your answers to Question 1. Submit your completed Excel spreadsheet.
Answer #1

(a)

Following table shows the group totals

G1 G2 G3 G4
3 10 3 20
2 0 4 11
6 7 5 8
4 8 2
4 0 14
7
2
Total 28 25 12 55

Following table shows the grand total and sum of squares of values:

G G^2
3 9
2 4
6 36
4 16
4 16
7 49
2 4
10 100
0 0
7 49
8 64
0 0
3 9
4 16
5 25
20 400
11 121
8 64
2 4
14 196
Total 120 1182

Now,

\sum G_{1}=165,\sum G_{2}=145,\sum G_{3}=140,n_{1}=5,n_{2}=5,n_{3}=5,\sum G =450,\sum G^{2}=13806
Now

Now

Now

Since there are 5 different groups so we have k=4. Therefore degree of freedoms are:

-------------

Now

F test statistics is

So p-value of the test using excel function "=FDIST(3.11,3,16)" is 0.0559. Since P-value is not less than 0.05 so there is evidence to conclude that average daily sales for four different dealerships in four separate states are equal.

(b)

Using excel Data Analysis - One fcator ANOVA :

Following is the output: